OpenStudy (anonymous):

The variables x and y vary directly. Use the given values to write an equation that relates x and y: x = 3, y = 15

OpenStudy (anonymous):

6x-3=y

OpenStudy (anonymous):

If they vary directly, that means y = cx for some value c. Solve for c by plugging in the x and y we know 15 = 3c so, c = 5 The equation that relates x and y is y = 5x

OpenStudy (the_fizicx99):

The formula for direct variation is y = kx 15 = 3x /3 /3 x = 5 So, 15 = 3(5) 15 = 15 so true :3

OpenStudy (anonymous):

it said that make an equation so i make it complicated but the simplest is 5x=y :D

OpenStudy (anonymous):

Thank you